The components of a force are Fx = -35 N and Fy = 49 N.

What angle does this force make with the positive x axis?

The components of an electric field are Ex = -0.2 N/C and Ey = 0.2 N/C.

What angle does this electric field make with the positive x axis?
Answer #1

Solution) Given components of force along X and Y direction

Along X direction

Fx = |F|cos(theeta) •••••(1)

Fy = |F|sin(theeta) ••••••••(2)

|F| = (Fx^2 + Fy^2)^(1/2)

Therefore on dividing (2) by (1) we get

|F|sin(theeta)/|F|cos(theeta) = Fy/Fx

Tan(theeta) = Fy/Fx

Theeta = tan inverse(Fy/Fx)

Same concept applies to electric field E also .